The refractive indices of the core and cladding materials of an optical fibre are 1.64  and 1.53 respectively. Calculate the numerical aperture and light gathering capacity  of the fibre.


1
Expert's answer
2021-06-09T08:01:22-0400

"NA=\sqrt{n_2^2-n_1^2}=\sqrt{1.64^2-1.53^2}=0.59,"

"\theta=\arcsin \frac{n_1}{n_2}=\arcsin\frac{1.53}{1.64}=68.9°."